You have a right triangle labeled RTS. T is the 90˚ angle. S equals 22˚. I have to find the hypotenuse (I labeled it x) of this right triangle. There is also a measurement given to me for side TS. It is labeled 20.
So here is what I know:
Side TS= 20˚
side TS= the adj side of my know angle ( Question: How do I find the hypotenuse of triangle RTS? What formula would I use to solve for the hypotenuse of triangle RTS?
I know I have to use cosine, tangent, or sine formulas.
Thank you and please get back to me soon.
*Rosella Black*

Answer by Okeke_Christian(26)

Rosella,
Thanks for the question.
You should use the Cosine Ratio:

Cosθ = adj/hyp
For ease of remembrance, use the mnemonic "SOH CAH TOA" (say /so-car-to-ah/)
Where,
SOH means Sineθ = opp/hyp
CAH means Cosθ = adj/hyp
TOA means Tanθ = opp/adj

I do hope you find this useful!